OpenStudy (anonymous):

Which of the following expressions would appear furthest to the right on a number line? √3+5 7 √36 15/2

OpenStudy (anonymous):

\[\sqrt{36}=6\] \[\frac{15}{2}=7.5\] and since the square root of three is less then two you know \[\sqrt{3}+5<7\] so answer is 7.5